OpenStudy (anonymous):

For questions 1–2, what is the graph of the system? y ≤ x + 4 2x + y ≤ –4

OpenStudy (anonymous):

help please!

OpenStudy (mathmate):

 The graph shows y=x, and y=x+4, meaning the graph of y=x+4 is just 4 units higher than y=x. Can you figure out which half-plane satisfies \( y \le x+4 \) ?

OpenStudy (anonymous):

ya i got it thanks

OpenStudy (mathmate):

yw! :)